OpenStudy (anonymous):

Which of the following represents the area of a rectangle whose length is 3x – 7 and whose width is x + 2? A. 3x^2 – 14 B. 3x^2 + x – 14 C. 3x^2 – x – 14 D. 3x^2 + 14

OpenStudy (anonymous):

A = L x W So, (3x-7)(x+2) =

OpenStudy (anonymous):

3X-14

OpenStudy (anonymous):

3X^2-14

OpenStudy (anonymous):

Missing a term....do you know FOIL?

OpenStudy (anonymous):

First 3x * x = 3x^2 Outer 3x * 2 = 6x Inner -7 * x = -7x Last -7 * 2 = 14 Add all those terms, you get 3x^2 (-7x +6x) +14 3x^2 -x +14

OpenStudy (anonymous):

that isn't 1 of the choices

OpenStudy (anonymous):

Damn...sorry...my sign was wrong in Last...should have been -14....

OpenStudy (anonymous):

I hope you see at least how I got there using the FOIL method

OpenStudy (anonymous):

so it's 3x^2 -x -14?

OpenStudy (anonymous):

yes